# Atlas: A Framework for ML Lifecycle Provenance & Transparency

### v1

@@ 5.1 Kubeflow Integration
5.2 TEE Service Claims
For each attestation request, the client generates a JSON Web Token  [53], which adds custom claims to represent Atlas-specific information including issuer identification and public key for signature verification, temporal validity through expiration and effective timestamps, hardware TEE information and software measurements, and golden values used in verification policy enforcement
Verification Service Implementation. The metadata sidecar also serves as a verification endpoint allowing pipeline components to validate artifact integrity against stored attestations. We optimize the performance of our staged verification system in three ways: 1) by processing changes incrementally and caching to avoid re-verifying unchanged components, 2) via batch processing of verification operations, and 3) parallel verification paths for independent component classes.
@@ 5.2 TEE Service Claims
For provenance chain validation, our implementation parallelizes verification of cross-component dependencies, version compatibility and evaluation result consistency. The system also maintains verification checkpoints that serve as trusted reference points, enabling partial verification from the last known good state instead of complete chain recomputation.
@@
5.3 Verification Flow
Figure 2: Verification workflow implementation showing how components are classified and processed: Changes trigger incremental validation that preserves cached states for unchanged components. The system groups related artifacts (training data, model weights, configurations) for batch processing, validates their relationships in parallel, and selectively invalidates only affected components on failures. This approach reduces verification overhead while maintaining security guarantees.
5.4 Framework Adaptability
@@
6 Evaluation
We validate our framework through a security analysis, preliminary performance testing and two case studies: BERT Meta [3, 4] and BGE reranker [5].
6.1 Security Analysis
Against hub threats, the Atlas verification system ensures artifact integrity through cryptographic measurements and signatures. The provenance chain maintains verifiable links between artifacts, preventing mismatched signatures or compromised dependencies from propagating.
To address artifact producer threats, Atlas relies on comprehensive provenance tracking that captures all dependencies and modifications. The transparency layer provides an immutable record of pipeline operations, enabling detection of undeclared dependencies or intentional omissions.
6.2 Preliminary Performance Analysis
Our implementation addresses the computational overhead of security mechanisms through several optimization strategies. Preliminary tests demonstrate that Atlas’ security mechanisms introduce minimal overhead, with training overhead remaining under 8% across all test cases. The verification processes scale linearly with model size, while our caching strategies (see §5) effectively optimize performance for iterative ML workflows; we observe a reduction of up to 50% of verification latency compared to the unoptimized implementation, and near-constant time for cached component verification.
For large-scale operations, parallel processing capabilities maintain performance at scale through: 1) concurrent verification operations when dependencies allow; 2) cache optimization to reduce repeated verification costs; 3) selective cache invalidation for error handling.
We plan to conduct more extensive performance benchmarking as part of future work.
6.3 Case Studies
BERT was selected for its complex architecture and widespread production use. Our implementation covers the complete lifecycle from pre-trained model fine-tuning through deployment, with isolated execution environments for attention computations and weight updates.
To validate Atlas’ adaptability, we also implement security controls for BGE reranker on Intel Xeon processors. Atlas secures both instruction-based and layerwise reranking configurations, with training data structured as JSON records containing query-positive-negative text triplets. The provenance system tracks model adaptations across both base model variations (Gemma-2B and MiniCPM-2B), maintaining verifiable records of hyperparameters and training progressions.
Both implementations demonstrate performance characteristics consistent with our analysis in  §6.2, while preserving training efficiency and security guarantees.
7 Case Study: BERT Pipeline
Our proof-of-concept implementation demonstrates the framework’s capabilities through securing a BERT model development and deployment pipeline. We chose BERT Meta [3, 4] for this case study due to its complex architecture, widespread use in production environments, and frequent fine-tuning requirements that test our framework’s adaptability.
The implementation covers the complete model lifecycle from pre-trained model adaptation through deployment. For BERT’s transformer architecture, we implemented isolated execution environments for attention computations and weight updates. The hardware-backed attestation system provides continuous runtime verification during fine-tuning operations, while our provenance system maintains verifiable records of all model evolution steps from the base model through task-specific adaptations.
Performance analysis shows practical viability with training overhead under 8%, primarily from manifest generation, and verification processes scaling linearly with model size. The system’s parallel processing capabilities and cache optimization strategies maintained performance at scale.
8 Discussion and Conclusion
Our implementation and evaluation reveal several key considerations for practical deployment ML pipeline security. Organizations must balance security requirements with operational efficiency, considering factors like verification frequency, attestation depth, and computational overhead. Atlas’ modular design enables customization of these tradeoffs while maintaining core security guarantees.
Current hardware security architectures present integration challenges, particularly around GPU and accelerator attestation support. While Atlas provides secure boundaries for CPU-GPU data transfers, emerging technologies like confidential GPU computing will enable more comprehensive coverage. Integration with legacy HPC systems also requires careful consideration, as traditional job schedulers often lack modern observability capabilities needed for fine-grained monitoring and attestation.
9 Adoption Considerations
9.1 Storage Optimizations
9.1.1 Storage and Scalability
@@
9.1.2 Decomposed Storage Model
@@
9.1.3 Hybrid Storage Architecture
@@
9.2 Deployment Guidelines
9.2.1 Organizational Adaptations
@@
9.2.2 Security Requirement Balance
@@
9.2.3 Computational Overhead Management
@@ 10 Future Work
Our work demonstrates that robust security measures can be integrated into ML lifecycles without significant performance impact. The combination of hardware-backed security guarantees with provenance tracking provides a foundation for securing ML pipelines. Our case studies show Atlas’ ability to maintain strong security properties while preserving established development practices.
We identify three directions for further enhancing ML pipeline security: 1. extending hardware security coverage to specialized ML accelerators will enable more comprehensive protection of ML computations; 2. standardizing provenance tracking across the industry will significantly improve interoperability and adoption; 3. developing robust algorithmic verification methods and model guardrails will strengthen protection against sophisticated attacks targeting model behavior, allowing for flexible integration of emerging defenses.

### v2

@@ B.1 Kubeflow Integration
B.2 C2PA Metadata Examples
A typical execution record captured by our system looks like:
        "execution": {
            "name": "training-run-132",
            "state": "RUNNING",
            "pipeline_spec": {
                "parameters": {
                    "learning_rate": 0.001,
                    "batch_size": 32,
                    "random_seed": 42,
                    "optimizer_config": {
                        "type": "Adam",
                        "beta1": 0.9,
                        "beta2": 0.999
                "runtime_config": {
                    "gcs_output_directory": "gs://...",
                    "tensorflow_version": "2.9.0"
For each execution, our system adds corresponding integrity measurements and verification records:
        "integrity_measurement": {
            "component_id": "training-run-132",
            "tdx_quote": "base64:...",
            "environment_hash": "sha256:...",
            "timestamp": "2024-01-15T10:30:00Z",
            "parameter_hash": "sha256:..."
B.3 ML System Monitoring Procedures
Our proof-of-concept implementation leverages several techniques to monitor ML pipeline activities with minimal intrusion into existing workflows. The implementation focuses on collecting runtime data about model weights, hyperparameters, and execution context while maintaining performance and compatibility with established ML frameworks.
B.3.1 File System Monitoring in the Atlas Framework
The Atlas sidecar implements a file system monitor written in Rust that detects checkpoint creation and modification events:
1:procedure InitCheckpointMonitor($dir,client$)
2:   Initialize directory and client references
3:   Create empty checksum tracking map
4:   Setup file system watcher
5:end procedure
6:procedure SetupWatcher
7:   Create event listener for file changes
8:   Start background monitoring thread
9:   Register directory for change notifications
10:end procedure
11:procedure ScanCheckpoints
12:   for each checkpoint file in directory do
13:      Compute file cryptographic checksum
14:      Store checksum in tracking map
15:      Register existing checkpoint in metadata
16:   end for
17:end procedure
18:procedure OnFileCreated($file$)
19:   if file is checkpoint type then
20:      Compute checksum and record creation
21:      Update checksum tracking map
22:   end if
23:end procedure
24:procedure OnFileModified($file$)
25:   if file is checkpoint type then
26:      Compute new checksum
27:      Retrieve old checksum from map
28:      if checksums differ then
29:         Update tracking map
30:         Record modification in metadata
31:      end if
32:   end if
33:end procedure
B.3.2 Callback/Hook Registration in PyTorch
For our BERT Meta case study, we implemented a callback system that integrates with PyTorch’s event mechanisms. The model monitoring component operates as follows:
1:procedure InitModelMonitor($model,client$)
2:   Store references to model and client
3:   Register monitoring hooks on model
4:end procedure
5:procedure RegisterHooks
6:   for each layer module in model do
7:      if module is neural network layer then
8:         Attach forward hook for activation capture
9:         if module has trainable weights then
10:            Attach gradient hook for updates
11:         end if
12:      end if
13:   end for
14:end procedure
15:procedure ForwardHook($module,input,output$)
16:   Calculate unique layer identifier
17:   Extract statistical metrics from output
18:   Record activation data to metadata store
19:end procedure
20:procedure GradientHook($gradient$)
21:   Calculate gradient magnitude
22:   Record gradient event with timestamp
23:end procedure
Additionally,for some cases we extended PyTorch’s standard training loop with epoch-level callbacks [82]:
1:procedure InitTrainingCallback($client$)
2:   Store reference to metadata client
3:end procedure
4:procedure OnEpochStart($epoch,optimizer$)
5:   Create hash of optimizer configuration
6:   Record epoch start event with config hash
7:end procedure
8:procedure OnEpochEnd($epoch,metrics,model$)
9:   Capture cryptographic model state snapshot
10:   Record completion with metrics and snapshot
11:end procedure
These hooks operate with minimal overhead while providing comprehensive visibility into the model’s evolution.
B.3.3 Configuration Wrappers in the Atlas Framework
To extract hyperparameter access and modifications from PyTorch, we implement transparent wrapper classes:
1:procedure InitTrackedConfig($config,client$)
2:   Store config and client references
3:   Initialize version counter to zero
4:   Record initial configuration state
5:end procedure
6:procedure Get($key$)
7:   Log access event to metadata store
8:   Return value for requested key
9:end procedure
10:procedure Set($key,value$)
11:   Retrieve current value for key
12:   Update configuration with new value
13:   Increment version counter
14:   Record modification in metadata
15:   Update configuration state hash
16:end procedure
17:procedure RecordState
18:   Generate hash of current configuration
19:   Store versioned snapshot in metadata
20:end procedure
21:procedure GetConfig
22:   Return copy of configuration
23:end procedure
B.4 Integration Between Framework and ML Pipeline
For the BERT Meta case study, we developed an integration layer that allows the framework components to interact with the Python ML pipeline:
1:procedure InitBridgeService($endpoint,dir$)
2:   Create metadata client connection
3:   Initialize checkpoint monitor
4:   Start bridge service
5:end procedure
6:procedure InitConfig($configJSON$)
7:   Parse configuration from JSON
8:   Create configuration tracking wrapper
9:   Generate unique tracking identifier
10:   return tracking identifier
11:end procedure
12:procedure ScanCheckpoints
13:   Scan and register existing checkpoints
14:end procedure
15:procedure RecordEvent($type,data$)
16:   if $type$ is epoch start then
17:      Record epoch initialization
18:   else if $type$ is epoch end then
19:      Record epoch completion with metrics
20:   else if $type$ is layer activation then
21:      Record layer output statistics
22:   else if $type$ is gradient event then
23:      Record gradient flow information
24:   end if
25:end procedure
On the Python side, we implement a complementary bridge client:
1:procedure InitBridgeClient($socketPath$)
2:   Connect to monitoring bridge service
3:end procedure
4:procedure SetupMonitoring($model,config,dir$)
5:   Serialize configuration to JSON
6:   Initialize configuration tracking
7:   Scan existing model checkpoints
8:   Create model monitoring hooks
9:   Setup training loop callbacks
10:   return monitoring components
11:end procedure
12:procedure RecordLayerData($layer,stats$)
13:   Package activation statistics
14:   Send to bridge as layer event
15:end procedure
16:procedure RecordGradient($magnitude,time$)
17:   Package gradient information
18:   Send to bridge as gradient event
19:end procedure
20:procedure RecordEpochStart($epoch,optHash$)
21:   Package epoch initialization data
22:   Send to bridge as epoch start event
23:end procedure
24:procedure RecordEpochEnd($epoch,metrics,hash$)
25:   Package completion metrics
26:   Send to bridge as epoch end event
27:end procedure
Figure 2: Altas sidecar collector showing data flow between Python ML environment and framework: The diagram illustrates how monitoring events are cached before being committed to the transparency log.
This hybrid architecture enables monitoring of the BERT training process with minimal modifications to the existing pipeline, while leveraging efficient system-level operations for the monitoring infrastructure.
B.5 Attestation Client Storage Optimization
As a storage optimization, the attestation client’s metadata sidecar first stores all generated C2PA manifests in a local cache layer before being committed to the transparency log. The local cache maintains an indexed hierarchy of manifests for efficient validation during pipeline execution, before final storage in Rekor for tamper-evident provenance tracking.
More specifically, we decompose manifests into constituent components within the cache. The C2PA metadata assertions, claim signatures, and pipeline metadata are stored separately, with relationships maintained through a reference system. This approach enables efficient updates to specific manifest components, reduced storage redundancy, optimized query performance, and scalable version tracking.
B.6 Verification Service Optimizations
Figure 3: Verification workflow implementation for Atlas components. The system classifies artifacts and processes changes incrementally, preserving cached states for unchanged components. Related artifacts are grouped for batch processing, with parallel validation of relationships. During errors, only affected components are invalidated, reducing verification overhead while maintaining security guarantees.
@@ B.6 Verification Service Optimizations
For provenance chain validation, our verification service implementation parallelizes verification of cross-component dependencies, version compatibility and evaluation result consistency. The system also maintains verification checkpoints that serve as trusted reference points, enabling partial verification from the last known good state instead of complete chain recomputation.
@@
B.7 Framework Adaptability
@@
Appendix C Adoption Considerations
C.1 Storage Optimizations
C.1.1 Storage and Scalability
@@
C.1.2 Decomposed Storage Model
@@
C.1.3 Hybrid Storage Architecture
@@
C.2 Deployment Guidelines
C.2.1 Organizational Adaptations
@@
C.2.2 Security Requirement Balance
@@
C.2.3 Computational Overhead Management
@@ Appendix D Future Work
Throughout this process, Atlas provides tamper-evident records through its transparency log and TEE-based attestations. The verification service requires access to both the model artifacts and the cryptographic measurements in the transparency log to confirm the integrity of the complete ML pipeline, ensuring that no unauthorized modifications occurred during the model’s lifecycle.